Thomas’ Calculus Early Transcendentals 13ed

Study guide for quiz 03

Quiz problems include both the lecture contents and homework problems.

1. Section 2.6: Study definition of the limits in p119, p125, p131 and problem 6, homework 04. Find an example for each and for practice proving with these definitions.

2. Section 3.2: Study the definition of derivative, one-sided derivatives, and relation between differentiability and continuity (Theorem 1 and proof).

3. Section 3.3: Study derivative product rule, derivative quotient rule and applications

such as

\frac{d^{n}}{dx^{n}} (f(x)g(x)) =?, \frac{d}{dx}(f_{1}(x)f_{2}(x) ⋯ f_{n}(x)) =?

and derivative of determinants, etc.

1 Brief answers to selected homework problems

Figure 1: Brief answers to selected problems in section 2.6, part 1 Figure 2: Brief answers to selected problems in section 2.6, part 2

3 Figure 3: Brief answers to selected problems in section 3.2

4 Figure 4: Brief answers to selected problems in section 3.3

5